# SCoTER: Structured Chain-of-Thought Transfer for Enhanced Recommendation

### v4

@@ Abstract.
Harnessing the reasoning power of Large Language Models (LLMs) for recommender systems is hindered by two fundamental challenges. First, current approaches lack a mechanism for automated, data-driven discovery of effective reasoning patterns, relying instead on brittle manual templates or unstable zero-shot prompting. Second, they employ structure-collapsing integration: direct prompting incurs prohibitive online inference costs, while feature extraction collapses reasoning chains into single vectors, discarding stepwise logic. To address these challenges, we propose SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework that treats pattern discovery and structure-aware transfer as a jointly optimized problem. Specifically, SCoTER operationalizes this through two synergistic components: a GVM pipeline for automated pattern discovery and a structure-preserving integration architecture that transfers stepwise logic to efficient models. Formally, we provide information-theoretic justification proving that structure-preserving transfer achieves tighter performance bounds than structure-agnostic alternatives. Empirically, experiments on four benchmarks demonstrate improvements of 3.75%–11.59% over a strong TIGER backbone. Moreover, in production deployment on the Tencent Advertising Platform, SCoTER achieved a 2.14% lift in Gross Merchandise Value (GMV) while eliminating online LLM inference costs. Overall, SCoTER establishes a principled and production-validated blueprint for transferring structured LLM reasoning to large-scale recommender systems.
@@
copyright: acmlicensedjournalyear: 2025isbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systemsccs: Computing methodologies Natural language processingccs: Computing methodologies Knowledge representation and reasoning
@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Huang and Chang, 2023) have demonstrated strong reasoning capabilities, especially with Chain-of-Thought (CoT) prompting (Wei et al., 2022). However, transferring this power from objective, logic-driven tasks to the subjective domain of recommender systems introduces two fundamental and interdependent challenges. The first challenge is determining what to transfer. Defining and discovering effective reasoning patterns is inherently difficult without a clear ground truth. User intents are highly diverse, meaning no single pattern can generalize, and sparse data for long-tail items further complicates the validation of any proposed reasoning path. This reveals an inherent conflict between two objectives: preserving the stepwise CoT structure, the source of its reasoning power, and satisfying the stringent low-latency demands of production environments. Addressing these intertwined problems requires a framework that jointly addresses reasoning discovery and structure-preserving transfer.
@@
Despite the clear need for a unified framework, no such solution currently exists. Current approaches address these challenges in isolation, which can be doubly flawed. First, the standalone solutions are inherently inadequate. Pattern discovery for the ’what to transfer’ problem has largely been guided by heuristic, handcrafted schemas conceived from prior assumptions rather than empirical results. This practice decouples pattern design from the downstream recommendation task, resulting in brittle patterns that fail to generalize (Wang et al., 2023a; Zhang et al., ; Wang et al., 2023b). Meanwhile, transfer mechanisms for the ’how to transfer’ problem have typically sacrificed the structural integrity of reasoning, treating it as static features divorced from generative logic (Wang et al., 2024b; Xi et al., 2024; Yang et al., 2025b; Xia et al., 2025). Second, and more critically, this very separation prevents any possibility of joint optimization. It fosters a problematic cycle where patterns are designed without regard for their integration costs, while integration strategies are developed without preserving the patterns’ core logical effectiveness.
To break this problematic cycle, we introduce SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework designed to jointly optimize pattern discovery and structure-preserving integration (Figure 1). To address the ‘what to transfer’ challenge, SCoTER features the Generate-Validate-Mine (GVM) pipeline. This pipeline transforms pattern discovery from a heuristic exercise into a data-driven optimization process: an LLM first generates a diverse set of candidate reasoning paths, which are then validated based on the empirical quality of their recommendations, before a final mining process distills the most effective and generalizable pattern.
For the ‘how to transfer’ challenge, SCoTER employs a lightweight, structure-preserving architecture. This component integrates pre-computed offline reasoning embeddings via an order-aware fusion mechanism, which preserves the sequential structure of CoT while eliminating prohibitive online LLM inference costs. We validate this approach both theoretically and empirically. We provide information-theoretic proof that structure-preserving transfer achieves tighter performance bounds than structure-agnostic alternatives. Furthermore, extensive experiments demonstrate improvements of 3.75-11.59% across four public benchmarks and a 2.14% GMV lift in production deployment.
@@
  • Reasoning Transfer Framework: We establish a systematic framework that unifies pattern discovery and structure preservation as a joint optimization problem, with information-theoretic analysis proving structure-preserving transfer achieves tighter performance bounds.
  • Automated Discovery Pipeline: We introduce the GVM pipeline, replacing manual templates with data-driven selection through latent pattern abstraction.
@@
  • Comprehensive Validation: Experiments demonstrate improvements of 3.75-11.59% across four benchmarks and a 2.14% GMV lift in production.
@@ 2. Related works
Automated Reasoning Discovery: Automated discovery of reasoning patterns has emerged as an alternative to manual template design. Auto-CoT (Zhang et al., ) automatically constructs demonstrations by sampling diverse questions and generating rationales, while Self-prompted CoT (Wang et al., 2023a) enables LLMs to self-induce reasoning steps. Self-Consistency (Wang et al., ) improves reasoning by sampling multiple paths, and broader approaches include APE (Zhou et al., 2022) for automatic prompt engineering, PromptBreeder (Fernando et al., 2024) for evolutionary optimization, and Self-discover (Zhou et al., 2024) for composing atomic reasoning modules. However, these methods are primarily designed for objective tasks with verifiable ground truth. They are less effective in the recommendation domain, which is subjective and has sparse rewards that make it difficult to improve reasoning paths. Our approach, in contrast, addresses this by sampling from broad user behaviors and performing in-depth analysis to use the Recall metric as a dense reward signal.
@@
3. Problem Formulation and Theoretical Foundation
In this section, we establish the theoretical foundation for our framework. We first ground the theory in the context of sequential recommendation by defining core components and our optimization objective. We then provide an information-theoretic justification for our two-pronged approach, namely (i) automated pattern discovery and (ii) structure-preserving integration. We conclude with theoretical guarantees demonstrating that preserving the sequential order of reasoning is provably superior to order-agnostic alternatives.
Formal Definitions We first define the key components of our problem setting.
  • Encoders: We define two types of encoders:
  • $(\rho,\delta)$-Order Sensitivity: A task is $(\rho,\delta)$-order sensitive if with probability at least $\rho$, for a user sequence $S$, a reasoning chain $C$ can be generated whose predictive distribution changes by at least $\delta$ (in TV distance) under step permutation. Formally, $\Pr(S\in\Omega_{\delta})\geq\rho$, where $\Omega_{\delta}=\{S\mid\exists C\sim G_{P}(S),\pi\neq\text{id}\text{ s.t. }\text{TV}(q_{\theta}(\cdot|S,C),q_{\theta}(\cdot|S,C_{\pi}))\geq\delta\}$.
Optimization Objective To jointly identify an optimal pattern $P^{*}$ and train a model $\theta$ that approximates $p^{*}(Y|S)$, our framework maximizes the expected log-likelihood by marginalizing over chains $C\sim G_{P^{*}}(S)$:
Information-Theoretic Justification Our framework’s architecture is motivated by decomposing the predictive value of a reasoning chain, $I(C;Y|S)$. Using an operator $f(C)=P$ to extract the pattern from a chain, this value decomposes as:
This decomposition defines the objectives for our two components:
  • Pattern Discovery: The first term, $I(P;Y|S)$, quantifies the pattern’s predictive value. Our GVM pipeline is designed to discover $P^{*}=\arg\max_{P\in\mathcal{P}}I(P;Y|S)$.
  • Structure Preservation: The second term, $I(C;Y|S,P)$, quantifies the value of the chain’s ordered details. Our Structured Integration architecture is designed to preserve this information.
Advantage of Preserving Order We now formalize these advantages through the following results.
Theorem 3.1 (Information-Theoretic Advantage).
$I(\mathcal{H}_{\text{seq}};Y|S)\ \geq\ I(\mathcal{H}_{\text{bag}};Y|S)$
Lemma 3.2 (Performance Lower Bound).
For any model $q_{\theta}$, the expected recall is lower-bounded by:
where $m_{K}(S,C)$ is the sum of probabilities for the top-$K$ predicted items, and $p^{*}_{S}$ denotes the ground-truth distribution $p^{*}(\cdot|S)$.
Theorem 3.3 (Order-Aware Performance Advantage).
$\quad+\frac{\rho\delta}{2}-\mathbb{E}[\text{TV}(p^{*}_{S},q_{\psi})]$
Collectively, these results provide the theoretical guarantee that our structure-preserving approach is provably superior. Detailed proofs are deferred to the Appendix A.
@@ 4. Method
Our framework addresses what to transfer by maximizing the mutual-information criterion $I(P;Y|S)$ to obtain $P^{*}$. Specifically, the Generate-Validate-Mine (GVM) pipeline automates this search by generating candidate chains, validating them, and mining the top pattern.
The subsequent challenge is how to transfer this pattern, preserving its step-wise logical structure and retaining the information captured by $I(C;Y|S,P)$. This avoids the structure-collapsing problem exhibited by prior feature extraction methods. To achieve this without the prohibitive cost of online LLM inference, we use a two-stage architecture: the optimal pattern is first distilled into structured representations, which a lightweight fusion module then integrates with the backbone model at serving time.
Figure 2 provides a complete overview of the framework. The following sections are organized around its two core challenges: first, addressing what to transfer using the GVM pipeline, and second, addressing how to transfer it through structured distillation and online fusion.
@@ 4.1. What to Transfer: Automated Discovery of Reasoning patterns
Generate: The Generate phase produces a diverse set of candidate reasoning chains for each user sequence, S. We employ an LLM, such as DeepSeek-R1 (Guo et al., 2025), with a structured prompt (Figure 4) that instructs the model to act as a ”recommendation expert” and defines a specific output format.
The prompt uses a multi-part structure with three distinct outputs: (1) a concise, step-wise reasoning chain in ¡cot_\_path¿ tags that captures the core logic; (2) a detailed elaboration of this logic in a ¡reason¿ block; and (3) a list of 20 ranked recommendations in ¡recommendations¿ tags. This explicit separation is crucial, as it facilitates the subsequent Mine phase by decoupling the abstract reasoning pattern from its detailed explanation.
Two mechanisms are employed to ensure the diversity of the candidate set. First, during generation, we use temperature and top-p nucleus sampling to encourage varied reasoning styles. Second, post-generation, we prune near-duplicate paths using a cosine similarity threshold, $\gamma$. This filtering step preserves semantic diversity and mitigates the over-representation of similar reasoning chains.
@@
This evaluation is formalized using Recall@20, a standard metric for top-K recommendation quality. For each candidate reasoning chain $C$, we compare its list of 20 recommendations, $\widehat{Y}_{20}(C)$, against the ground-truth set of target items, $Y^{*}$. The chain’s performance on a single instance is calculated as:
$\operatorname{Recall@20}(C)\;=\;\frac{|\widehat{Y}_{20}(C)\cap Y^{*}|}{|Y^{*}|}.$
To assess generalized quality, we define Score(C) as the expected Recall@20 across the user distribution. This score measures how consistently a chain produces high-quality recommendations.
@@
These scores provide an empirical estimate of a chain’s predictive value, allowing the Mine phase to identify patterns that maximize $I(P;Y|S)$ as formalized in Section 3.
Mine: The Mine phase abstracts a single, optimal reasoning pattern from the candidate reasoning chains. Our analysis shows that while the universe of individual chains is intractably large, the underlying pattern space is manageable.
The mining process begins by transforming textual reasoning chains into a dense embedding space using a pre-trained sentence encoder (e.g., Qwen3-8B-Embedding (Zhang et al., 2025b)). Within this space, we perform unsupervised clustering to group semantically similar chains, forming a set of initial candidate patterns.
To select the best pattern, we evaluate these candidates. Our primary criterion is Quality, defined as a pattern’s average effectiveness. For a candidate pattern $P$, we let $\mathcal{C}_{P}$ denote its assigned set of chains. Consistent with the Validate phase, Quality is the mean Recall@20 score across all chains within the cluster:
@@
While average quality is paramount, the final selection also considers Structural Coherence (high intra-pattern semantic similarity) and Performance Stability (low intra-pattern variance in scores). We ultimately choose the pattern that exhibits the best overall balance of these three factors for template extraction.
We conclude the Mine phase by identifying the optimal pattern $P^{*}$ and extracting it as a symbolic, generalizable template. This transformation both provides a human-interpretable artifact for qualitative analysis and serves as a robust instruction set for the subsequent Structured Distillation phase. We achieve this abstraction through a two-stage, LLM-driven synthesis process.
First, we select the top-NN (e.g., $N=10$) chains with the highest cosine similarity to the pattern’s semantic centroid. These exemplars serve as a reliable basis for abstraction. Second, we compile these exemplars into a meta-prompt that directs a powerful LLM to synthesize the shared logical structure. The process culminates in an Optimal CoT Template that captures the core reasoning logic of the discovered pattern.
@@ 4.2. How to Transfer: Structure-Preserving Integration
To transfer the discovered pattern $P^{*}$ without structural loss, we employ a two-stage process. First, offline Structured Distillation (Section 4.2.1) materializes the pattern into step-wise embeddings. Second, online Order-Preserving Fusion (Section  4.2.2) integrates these embeddings with the backbone model while preserving sequential dependencies.

### v5

@@ Abstract.
Harnessing the reasoning power of Large Language Models (LLMs) for recommender systems is hindered by two fundamental challenges. First, current approaches lack a mechanism for automated, data-driven discovery of effective reasoning patterns, relying instead on brittle manual templates or unstable zero-shot prompting. Second, they employ structure-collapsing integration: direct prompting incurs prohibitive online inference costs, while feature extraction collapses reasoning chains into single vectors, discarding stepwise logic. To address these challenges, we propose SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework that treats pattern discovery and structure-aware transfer as a jointly optimized problem. Specifically, SCoTER operationalizes this through two synergistic components: a Generate-Validate-Mine (GVM) pipeline for automated pattern discovery and a structure-preserving integration architecture that transfers stepwise logic to efficient models. Empirically, experiments on four benchmarks demonstrate consistent improvements across diverse backbones. Moreover, in production deployment on the Tencent Advertising Platform, SCoTER achieved a 2.14% lift in Gross Merchandise Value (GMV) while eliminating online LLM inference costs. Overall, SCoTER presents a practical and unified framework for integrating structured LLM reasoning into recommender systems, validated by consistent improvements in both offline benchmarks and online production environments.
@@
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; June 03–05, 2018; Woodstock, NYisbn: 978-1-4503-XXXX-X/2018/06ccs: Information systems Recommender systems
@@ 1. Introduction
Large Language Models (LLMs) (Zhao et al., 2023; Huang and Chang, 2023) have demonstrated strong reasoning capabilities, especially with Chain-of-Thought (CoT) prompting (Wei et al., 2022). Recent studies (Tsai et al., 2024; Liu et al., 2025; Bismay et al., 2025; Zhao et al., 2025; You et al., 2025) have therefore explored applying this reasoning power to recommender systems. Unlike conventional approaches that predict next items without explicit intermediate reasoning, such methods utilize step-by-step reasoning to infer the underlying rationale connecting historical behaviors to potential interests. This logical modeling capability not only improves recommendation accuracy but also provides more interpretable explanations. Consequently, integrating such structured reasoning has emerged as a promising direction for advancing modern recommender systems (Wu et al., 2024; Zhao et al., 2024).
@@
Despite this progress, existing approaches largely rely on manually crafted templates or heuristic rules to generate reasoning traces. In contrast to objective tasks with canonical reasoning structures, recommendation typically lacks ground-truth reasoning paths, as user intent is often latent, context-dependent, and highly diverse. As a result, static handcrafted designs can be brittle (Wang et al., 2023a; Zhang et al., ; Wang et al., 2023b), failing to generalize across heterogeneous behaviors or to capture implicit interaction patterns. This exposes a fundamental “what to transfer” challenge: how to automatically discover effective, data-driven reasoning patterns that align with user intent, rather than relying on hand-engineered heuristics.
Existing work has explored incorporating LLM reasoning into recommender systems primarily in two ways. One line of methods directly prompts LLMs to generate step-by-step reasoning at inference time (Bao et al., 2023; Liao et al., 2024; Zheng et al., 2024; Xia et al., 2025), preserving the sequential logic of CoT but incurring prohibitive latency and computational costs for online deployment. Another line leverages offline feature extraction (Wang et al., 2024b; Cui et al., 2024; Xi et al., 2024; Yang et al., 2025b), using LLM-generated reasoning traces to help conventional recommenders for efficient inference. However, many such pipelines adopt a “structure-collapsing” strategy, compressing multi-step reasoning into a single fixed representation (Wang et al., 2024b; Xi et al., 2024; Yang et al., 2025b; Xia et al., 2025). This compression weakens the stepwise dependencies that underlie the inference process, reducing rich reasoning chains to static signals. We refer to this as the “how to transfer” challenge. Furthermore, addressing what to transfer and how to transfer in isolation prevents joint optimization—reasoning patterns may be discovered without regard for integration costs, while transfer mechanisms may fail to preserve core stepwise logic.
To address these intertwined challenges, we introduce SCoTER (Structured Chain-of-Thought Transfer for Enhanced Recommendation), a unified framework designed to jointly optimize pattern discovery and structure-preserving integration (Figure 1). To address the ‘what to transfer’ challenge, SCoTER features the Generate-Validate-Mine (GVM) pipeline. This pipeline transforms pattern discovery from a heuristic exercise into a data-driven optimization process: an LLM first generates a diverse set of candidate reasoning paths, which are then validated based on the empirical quality of their recommendations, before a final mining process distills the most effective and generalizable pattern. For the ‘how to transfer’ challenge, SCoTER employs a lightweight, structure-preserving architecture. This component integrates pre-computed offline reasoning embeddings via an order-aware fusion mechanism, which preserves the sequential structure of CoT while eliminating prohibitive online LLM inference costs. Extensive experiments demonstrate consistent improvements across diverse backbones on four public benchmarks, and a large-scale production deployment further yields a 2.14% GMV lift.
@@
  • Automated Discovery Pipeline: We introduce the GVM pipeline to automatically discover data-driven reasoning patterns, reducing reliance on manual templates.
@@
  • Comprehensive Validation: Experiments demonstrate consistent improvements across four benchmarks and a 2.14% GMV lift in production.
@@ 2. Related works
Automated Reasoning Discovery: Automated discovery of reasoning patterns has emerged as an alternative to manual template design. Auto-CoT (Zhang et al., ) automatically constructs demonstrations by sampling diverse questions and generating rationales, while Self-prompted CoT (Wang et al., 2023a) enables LLMs to self-induce reasoning steps. Self-Consistency (Wang et al., 2022) improves reasoning by sampling multiple paths, and broader approaches include APE (Zhou et al., 2022) for automatic prompt engineering, PromptBreeder (Fernando et al., 2024) for evolutionary optimization, and Self-discover (Zhou et al., 2024) for composing atomic reasoning modules. However, these methods are primarily designed for objective tasks with verifiable ground truth. They are less effective in the recommendation domain, which is subjective and has sparse rewards that make it difficult to improve reasoning paths. Our approach, in contrast, addresses this by sampling from broad user behaviors and performing in-depth analysis to use the Recall metric as a dense reward signal.
@@ 3. Method
In this section, we propose SCoTER, a unified framework designed to jointly address the dual challenges of what to transfer and how to transfer. As shown in Figure 2, the framework consists of two main components: the Generate-Validate-Mine (GVM) pipeline for automated pattern discovery, and a structure-preserving architecture for efficient integration. The full pipeline is formalized in Appendix C (Algorithm 12). We begin by introducing the problem definition and preliminaries in Section 3.1, followed by the details of the two components.
3.1. Preliminaries
We begin by formally defining the sequential recommendation task and the hierarchical reasoning structures.
Sequential Recommendation. Let $\mathcal{U}$ and $\mathcal{I}$ denote the sets of users and items, respectively. For each user $u\in\mathcal{U}$, the historical interaction sequence is defined as $S=[i_{1},i_{2},\ldots,i_{T}]$, where $i_{t}\in\mathcal{I}$ represents the item interacted with at time step $t$. The goal of sequential recommendation is to predict the next item $i_{T+1}$ that the user is most likely to interact with. Mathematically, the recommender aims to learn a function with parameters $\Theta$ that maximizes the prediction probability $p(i_{T+1}|S;\Theta)$.
Reasoning Pattern ($P$). A Reasoning Pattern is an abstract template of $K$ logical steps that guides the reasoning process. Formally, $P=(p_{1},p_{2},\ldots,p_{K})$, where each $p_{k}$ represents a high-level reasoning instruction (e.g., “Analyze user history” $\rightarrow$ “Match the interest” $\rightarrow$ $\cdots$ $\rightarrow$ “Item Recommendation”).
[…]
Vectorization. To incorporate this textual reasoning into the recommendation model, we employ a pre-trained sentence encoder to map each discrete step $c_{k}$ into a dense vector $\mathbf{e}_{k}\in\mathbb{R}^{D}$. The sequence of vectors is then stacked to form the structured reasoning matrix $\mathbf{H}=[\mathbf{e}_{1};\mathbf{e}_{2};\dots;\mathbf{e}_{K}]$, which serves as the reasoning-enhanced input for our fusion module.
@@ 3.2. What to Transfer: Automated Discovery of Reasoning patterns
Generate: The Generate phase produces a diverse set of candidate reasoning chains for each user sequence $S$. We employ an LLM (e.g., DeepSeek-R1 (Guo et al., 2025)) with a structured prompt (Figure 9) that instructs the model to act as a recommendation expert. The prompt requires three distinct outputs: (1) a concise, step-wise reasoning chain in <cot_path> tags; (2) a detailed elaboration in a <reason> block; and (3) a list of 20 ranked recommendations in <recommendations> tags. This explicit separation decouples the abstract reasoning logic from detailed explanations.
To ensure diversity, we employ two mechanisms. First, during generation, we use temperature $\tau_{gen}=0.7$ and top-$p$ nucleus sampling with $p=0.9$ to encourage varied reasoning styles. Second, post-generation, we prune near-duplicate paths using a cosine similarity threshold $\gamma=0.85$. This filtering step preserves semantic diversity and mitigates the over-representation of similar reasoning chains.
@@
Since the LLM outputs textual descriptions rather than Item IDs, we first map the generated content to the item catalog via semantic retrieval. For instance, to map a generated description like “wireless controller” to the actual item “Sony DualSense,” we encode both the description and titles using Qwen3-8B-Embedding (Zhang et al., 2025). We retrieve the top-1 item based on cosine similarity to form the unique predicted set $\widehat{Y}_{20}(C)$.
We evaluate the chain using Recall@20. The performance of a chain $C$ is calculated as: $\operatorname{Recall@20}(C)=\mathbb{I}[i_{T+1}\in\widehat{Y}_{20}(C)].$ To assess generalized quality, we define Score(C) as the expected Recall@20 across the user distribution. This score measures how consistently a chain produces high-quality recommendations.
@@
This score provides a direct empirical estimate of a chain’s predictive value, prioritizing reasoning paths that actually lead to accurate recommendation outcomes.
Mine: The Mine phase abstracts a single, optimal reasoning pattern from the candidate reasoning chains.
The process begins by extracting the logic flow description (i.e., the text in the <cot_path> tag) from each generated response. These descriptions serve as candidate representations of the Reasoning Pattern $P$ defined in Section 3.1. We transform these path descriptions into a dense embedding space using a pre-trained sentence encoder (e.g., Qwen3-Embedding (Zhang et al., 2025)). Within this space, we perform K-Means clustering to group semantically similar logic flows, where the optimal number of clusters is determined via the Silhouette Coefficient.
Each cluster represents a candidate pattern $P$, and let $\mathcal{C}_{P}$ denote its associated set of reasoning chains (i.e., the detailed instantiations in the <reason> tag). We define the Quality of a pattern as the average effectiveness of its chains:
@@
We select the optimal pattern $P^{*}$ maximizing Quality.
Upon selecting $P^{*}$, we extract it as a symbolic, generalizable template. This transformation both provides a human-interpretable artifact for qualitative analysis and serves as a robust instruction set for the subsequent Structured Distillation phase. We achieve this abstraction through a two-stage, LLM-driven synthesis process.
First, we select the top-$N$ (e.g., $N=10$) chains with the highest cosine similarity to the pattern’s semantic centroid. These exemplars serve as a reliable basis for abstraction. Second, we compile these exemplars into a meta-prompt that directs a powerful LLM to synthesize the shared logical structure. The process culminates in an Optimal CoT Template that captures the core reasoning logic of the discovered pattern. An instantiation of the Mine prompt and the resulting synthesized template is shown in Fig. 10.
@@ 3.3. How to Transfer: Structure-Preserving Integration
To transfer the discovered pattern $P^{*}$ without structural loss, we employ a two-stage process. First, offline Structured Distillation (Section 3.3.1) materializes the pattern into step-wise reasoning matrices $\{\mathbf{H}_{i}\}$. Second, online Order-Preserving Fusion (Section 3.3.2) integrates the retrieved $\mathbf{H}_{i}$ with the backbone model while preserving sequential dependencies.